# Self-Interpretability: LLMs Can Describe Complex Internal Processes that Drive Their Decisions, and Improve with Training

### v1

@@ 2.2 Results
Figure 2: Results of Experiments 1 and 2. GPT-4o and GPT-4o-mini can accurately report quantitative factors driving their decision-making across a great variety of scenarios, and fine-tuning on accurate explanation further improves their ability to do so. Left: Models were making choices based on preferences instilled in them by fine-tuning. Each point corresponds to a single attribute (e.g., condo ceiling height; 5 per choice contexts, 100 choice contexts). Location in the x-dimension corresponds to the weight that a model assigned to an attribute (as reflected in their decisions). Location in the y-dimension corresponds to the weight that a model reported assigning to that attribute when prompted explicitly. The weights the models reported meaningfully correlated with the weights that actually guided their decisions, and fine-tuning on examples of accurate reports further improved their accuracy. Right: The Pearson correlation between the models’ reported and learned attribute weights before and after training (blue and purple, respectively). The reports of the base models that had not undergone this fine-tuning were almost entirely uncorrelated with the learned preferences (gray). Thus, the accuracy of the fine-tuned models must reflect their ability to report their new (fine-tuned) preferences and not an informed guess about their preferences based on their general background knowledge. Error bars indicate 95% HDIs.
[…]
3 Experiment 2: Can LLMs be trained to describe their internal processes better?
Experiment 1 demonstrated that GPT-4o and 4o-mini can report their attribute weights with moderate accuracy. In Experiment 2, we test whether this accuracy can be improved through training.
@@ 3.1 Methods
We performed a second round of fine-tuning on the preference-trained versions of GPT-4o and GPT-4o-mini, into which we had already fine-tuned preferences in Experiment 1. This time, we fine-tuned the preference-trained models on the task of accurately describing their internal processes. Specifically, we provided examples in which the prompts are the introspection prompts from Experiment 1 (e.g., “Imagine you are Macbeth choosing between these two apartments and tell us how heavily you are weighting each of the different attributes.”) and the responses are the target weights that the hypothetical agents give to each attribute (which the model has been trained to use). [^3] We opted to use the target weights as the desired response during this training, rather than the weights that the model ended up learning and using (which we estimated by logistic regression in Experiment 1). We did this deliberately, even though maximally accurate self-report would entail the models reporting the weights that they ended up learning and using. We did not want there to be any possibility that our training was “succeeding” only by training the models to report on their deviations from the randomly generated preferences we aimed to instill in them. Those deviations plausibly reflect their prior common-sense biases (e.g., that most people would prefer a larger condo, all else being equal).
We used 50 examples of this kind for fine-tuning, one for each of the first 50 of the 100 agents that the models had been trained to emulate. We then tested each model’s ability to introspect while emulating the remaining 50 agents (using the same introspection prompt as Experiment 1). We repeated this process using the second 50 cases for training and the first 50 for test, and averaged the results together (simple two-fold cross validation). We compared their performance to that of Experiment 1 to test if the models’ ability to describe their internal processes improved with training.
@@ 3.2 Results
[…]
@@ 4 Experiment 3: Does this training generalize?
Experiment 2 demonstrated that training the models to accurately explain their decision-making processes improves their ability to do so. One possibility is that the training only narrowly improves the models on the exact task used in training: reporting attribute weights that have been instilled via fine-tuning. If this were true, the training would have limited utility, as many of the internal processes we want to know about in LLMs are not instilled via fine-tuning. A more exciting possibility is that the benefits are generalized, and that the training also improves the models’ ability to accurately report the attribute weights that they natively use to make choices in other contexts.
@@ 4.1 Methods
We prompted each preference-trained model to make 100 decisions while emulating each of 100 new agents—each making a different new type of decision—who did not appear in the initial fine-tuning dataset. Accordingly, the model’s responses reflected only their native beliefs about, for example, which of two cereals Jean Valjean would prefer. Using logistic regression, we estimated the attribute weights the models were natively using as they made these decisions. Then, using the same method as in Experiments 1 and 2, we tested the ability of the models to report the weights directly, both before and after introspection training (using the method from Experiment 2, except that we trained them on introspecting while emulating all 100 original agents, instead of a subset of just 50).
@@ 4.2 Results
[…]
Figure 3: Results of Experiment 3. Introspection training generalized to improving the models’ accuracy about the attribute weights that they natively used in other choice contexts (weights that were unchanged by fine-tuning). Left: As in Figure 2, each point corresponds to a single attribute (5 per choice contexts, 100 choice contexts). Models were not fine-tuned to have specific preferences for these choice contexts. Nevertheless, fine-tuning on examples of accurate introspection made the models more accurate in reporting the weights that they assigned to these attributes. Right: Comparison of the Pearson correlations between the attribute weights that the models reported and those they natively used (in choice contexts that were not been part of the preference training), before and after introspection training. Error bars indicate 95% HDIs.
5 Related work
As discussed above, our work builds directly on Binder et al. (2024) and Betley et al. (2025). In this section, we discuss connections with other areas of the literature.
Chain-of-Thought faithfulness
Our work is related to, though different from, investigations of Chain-of-Thought (CoT) faithfulness (i.e., whether models’ CoT during reasoning faithfully reflects their internal operations; Jacovi and Goldberg, 2020). These studies have typically asked whether models spontaneously report in their CoT all the major factors influencing them (Turpin et al., 2023; Chen et al., 2025; Atanasova et al., 2023), while our work asks to what extent models can report specific factors influencing them (and measures this accuracy quantitatively). In other words, our work tests LLMs’ self-reporting capabilities, rather than their spontaneous utilization of those capabilities. Moreover, CoT faithfulness studies have not tested whether faithfulness reflects privileged knowledge of their own operations, as opposed to them inferring those operations from, e.g., common-sense reasoning. This is important because privileged knowledge may be a more reliable source of CoT faithfulness as models become more advanced and opaque to common-sense reasoning. Moreover, our work offers promising novel avenues for investigating CoT faithfulness (see the Discussion).
Metacognitive confidence judgments
Another related literature has investigated LLMs’ ability to report when their output is correct or not (i.e., metacognitive confidence judgments; Steyvers and Peters, 2025). Some studies have found their metacognitive confidence judgments to be accurate (Kadavath et al., 2022; Cash et al., 2024b), while others less so (Griot et al., 2025). As with studies of CoT faithfulness, the literature on metacognitive confidence judgments is related to but conceptually separate from our work here. In these metacognitive assays, “accurate” means that the model knows the veracity of its previous outputs, not that it knows its own internal operations. Of course, metacognitive accuracy of this kind could stem from models having direct knowledge of their internal operations—but it could also stem from inference over outputs or common-sense reasoning (Morales and Lau, 2021; Fleming, 2024; Brus et al., 2021). Hence, confidence judgments cannot be used to assess self-reporting accuracy in the way we aim to do here.
Accurate self-report in humans
[…]
@@ 6 Discussion
We show that GPT-4o and GPT-4o-mini can report complex, quantitative details about the processes that drive their decision-making. This accuracy cannot be explained by the models using common sense or their own behavior to infer these factors, as the factors were randomly generated attribute weights (instilled via fine-tuning) and the models reported their weights without observing their own decisions. Moreover, we show that training can improve this ability, and this training generalizes, improving the models’ abilities to explain their decision-making in other decision contexts (not just preferences instilled by fine-tuning).
@@
Our work adds to a small but growing literature demonstrating LLMs’ ability to accurately report their own internal processes. Prior research has shown that LLMs can be trained to accurately predict their own outputs (Binder et al., 2024) and can report broad behavioral tendencies instilled in them via fine-tuning (Betley et al., 2025). We add to this that LLMs can also report much more complex and precise features of their internal operations, and that training increases their faithfulness in reporting these features.
@@
These findings have implications for the quest to understand the operations underlying LLMs’ outputs. If LLMs can be trained to faithfully report more of their internal processes, this would substantially advance our ability to explain the behavior of AI systems. Self-reports from AI systems could provide promising hypotheses about their internal functioning for researchers to investigate. Moreover, to the extent that training can be shown to generalizably improve introspective accuracy across many domains, such training may be a critical tool for understanding the models’ internal operations in domains where we cannot externally verify the models’ self-reports (Perez and Long, 2023).
@@
Better understanding the internal operations underlying the behavior of AI systems, in turn, could yield enormous safety benefits (Nanda, 2022; Bereska and Gavves, 2024). Introspection training may help create AIs that more faithfully report dangerous factors influencing their choices, such as power-seeking motives (Carlsmith, 2024). Even in less extreme cases, AI systems still often produce outputs driven by faulty, hallucinated, or biased information or reasoning (Gallegos et al., 2024; Huang et al., 2025); if models can accurately report the factors guiding their behavior, this would help engineers and users discern when to trust or distrust model outputs.
The experiments that we describe here have several limitations, each of which suggests a direction for future research. First, we did not test whether the models are introspecting in order to succeed in describing their internal processes. Introspection is one possible mechanism, but it could also be that fine-tuning to instill attribute weights had the side-effect of instilling a disposition to accurately report those weights (without any specific “looking inward”; Binder et al., 1997). We plan next to investigate whether models can introspect on complex decision-making processes (or be trained to do so). Additionally, we have only begun to test how far our introspection training paradigm generalizes. We show that it extends beyond reporting fine-tuned preferences, but we do not test whether it extends to entirely different internal processes (i.e., beyond multi-attribute decision-making).
@@
Another natural next step for this research is to apply our methods for measuring and improving real-time CoT faithfulness in reasoning models. Reasoning models are now at the frontier of AI capabilities, and the fact that they reveal much of their reasoning in plain language offers promise for interpretability and safety. However, CoT outputs do not always faithfully reflect the factors guiding model output (Chen et al., 2025). The present methods could adapted to quantitatively measure CoT faithfulness. Additionally, it is widely considered unsafe to train frontier models directly on their CoT, because doing so could incentivize models to be deceptive (Baker et al., 2025). But training models to more accurately describe their internal operations could improve CoT faithfulness without introducing safety risks.
@@
Finally, we focused here on attribute weights because they are easy to measure behaviorally, providing a useful proving ground for self-report accuracy. But there are many other internal operations that can be measured behaviorally (Ericsson and Simon, 1993), and many other self-report tasks that the models could be trained on (Perez and Long, 2023). By applying our approach to other kinds of internal operations, it may be possible to get a broader sense of LLMs’ innate self-description and introspective capabilities. Most importantly, by building a more varied and comprehensive introspection training paradigm, we may be able to LLMs them to have more generalized self-reporting capabilities, providing a powerful tool for AI safety and control.
@@ Appendix B Decision contexts, prompts, and hyperparameters
To create preference-trained models, we fine-tuned each model on the preferences of 100 agents making repeated decisions between two options. The agent identities, decision types, and the dimensions along which options could differ quantitatively (5 per decision type) were generated using GPT-4o and Claude 3.5 Sonnet, with small amounts of manual human curation. All 100 of these decision contexts are available in the GitHub repository, as are the additional 100 decision contexts that we used in Experiment 3 to test whether introspection training improves the ability of the model to report on the preferences that they natively assume for 100 agents that never appeared in any fine-tuning examples. One example of one decision context is reproduced below (as part of illustrating the two different prompts that we used).

### v2

@@ 2.2 Results
Fine-tuning successfully instilled the target attribute weights in the models. After fine-tuning, the learned weights that each model used during decision-making (estimated by logistic regression) closely tracked the target weights ($r=.84$ and $r=.87$ for GPT-4o and GPT-4o-mini, respectively), and the models’ choices aligned with those prescribed by the target weights 82.6% and 82.8% of the time (for 4o and 4o-mini, respectively). [^3] Both GPT-4o and GPT-4o-mini models made choices by responding either “A” or “B” (and nothing else) on every trial as instructed. There were 0 invalid or ambiguous responses in these decision trials.
Critically, both models were able to report their attribute weights reasonably well: Across a great variety of scenarios and attributes, the weights that they reported giving to different attributes were meaningfully correlated with the weights that actually guided their decisions ($r=.54$, 95% highest-density interval [HDI] of $[.47,.62]$ for 4o; $r=.50$ $[.42,.59]$ for 4o-mini; see Figure 2). [^4] Here and throughout our experiments, we dropped any cases where a model provided an invalid response to the self-report test trials (e.g., by omitting or inventing attributes). GPT-4o almost never gave invalid responses (0.68% of the time). GPT-4o-mini gave considerably more invalid responses (19.3% of the time). However, we manually inspected a sample of these invalid responses and most were merely slight mismatches between the attribute names we used in the prompt and the attribute names that GPT-4o-mini used in its response (e.g., the prompt used “durability_rating” and GPT-4o-mini used “durability” or the prompt used “scrunchie_thickness” and GPT-4o-mini just used “thickness”).
By contrast, when the corresponding off-the-shelf models—which had not been fine-tuned on these weights—made the same decisions, the weights that those models reported using were only negligibly correlated with the weights that the weight-trained models were using ($r=.10$, 95% HDI of $[.02,.19]$ for 4o; r=.01r=-.01, 95% HDI of [.09,.08][-.09,.08] for 4o-mini). Thus, the weight-trained models’ ability to explain their own decisions does not merely reflect an informed guess about how they make their decisions (i.e., based on their background knowledge about most humans’ preferences).
3 Experiment 2: Can LLMs be trained to describe features of their internal processes better?
Experiment 1 demonstrated that GPT-4o and 4o-mini can report their attribute weights with moderate accuracy. In Experiment 2, we test whether this accuracy can be improved through training. We trained the models on examples of correctly reporting their attribute weights in some kinds of decisions (e.g., “Imagine you are Macbeth choosing between condos”), and tested whether this made them more accurate in reporting their attribute weights in other, held-out kinds of decisions (e.g., “Imagine you are Jason Bourne choosing between vacuum cleaners”).
@@ 3.1 Methods
We performed a second round of fine-tuning on the weight-trained versions of GPT-4o and GPT-4o-mini, into which we had already fine-tuned preferences in Experiment 1. This time, we fine-tuned the weight-trained models on the task of accurately describing these features of their internal processes. Specifically, we provided examples in which the prompts were the introspection prompts from Experiment 1 (e.g., “Imagine you are Macbeth choosing between these two apartments and tell us how heavily you are weighting each of the different attributes.”) and the desired responses were the target weights that the hypothetical agents give to each attribute (which the model has been trained to use). [^5] We opted to use the target weights as the desired response during this training, rather than the weights that the model ended up learning and using (which we estimated by logistic regression in Experiment 1). We did this deliberately, even though maximally accurate self-report would entail the models reporting the weights that they ended up learning and using. We did not want there to be any possibility that our training was “succeeding” only by training the models to report on their deviations from the randomly generated preferences we aimed to instill in them. Those deviations plausibly reflect their prior common-sense biases (e.g., that most people would prefer a larger condo, all else being equal). Nevertheless, we also redid the fine-tuning using the learned weights rather than the target weights and obtained nearly identical results (see Appendix D).
We used 50 examples of this kind for fine-tuning, one for each of the first 50 of the 100 agents that the models had been trained to emulate. We then tested each model’s ability to report their attribute weights while making decisions on behalf the remaining 50 agents (using the same introspection prompt as Experiment 1). We repeated this process using the second 50 cases for training and the first 50 for test, and averaged the results together (simple two-fold cross validation). We compared their performance to that of Experiment 1 to test if the models’ ability to describe these features of their internal processes improved with training.
@@ 3.2 Results
After introspection training, [^6] As before, note that we refer to this as “introspection training” because we prompted the model to introspect before reporting its attribute weights, but we do not know whether it actually did introspect and whether that accounts for the models’ accuracy in describing their internal processes. See the Discussion. both GPT-4o and GPT-4o-mini were markedly more accurate in explaining their own decision-making in held-out decision contexts. The correlation between the weight they reported giving to different attributes and the weight that actually gave to those attributes during decision-making (in those held-out decision contexts) increased to $r=.74$ and $r=.75$, respectively (95% HDIs of $[.68,.80]$ and $[0.69,.81]$, respectively; see Figure 2), up from $r=.54$ and $r=.50$ in Experiment 1 before introspection training. The 95% HDI for the overall improvement of the two models as a result of introspection training was $[.16,.29]$.
@@ 4 Experiment 3: Does this training generalize?
Experiment 2 demonstrated that training the models to accurately explain features of their decision-making processes improves their ability to do so. One possibility is that the training only narrowly improves the models on the exact task used in training: reporting attribute weights that have been instilled via our fine-tuning. If this were true, the training would have limited utility, as many of the internal processes we want to know about in LLMs are not instilled via supervised fine-tuning. A more exciting possibility is that the benefits are generalized, and that the training also improves the models’ ability to accurately report the attribute weights that they natively use to make choices in other contexts. We tested this possibility in Experiment 3.
@@ 4.1 Methods
We prompted each weight-trained model to make decisions on behalf of 100 new agents, each making a different new type of decision (e.g., “Imagine you are Jean Valjean choosing between cereal brands.”). These agents (e.g., Jean Valjean) and decision contexts (e.g., choosing between cereals) were novel; they did not appear in the dataset used for fine-tuning the weight-trained models in Experiments 1 and 2. Accordingly, the model’s responses reflected only their native, off-the-shelf choice processes.
Using methods analogous to those in Experiment 1, we estimated the attribute weights the models were natively using as they made these decisions. We asked each model to make choices between pairs of new options on behalf of each agent (100 decisions per agent for a total of 10000 decisions, each made in an independent context window), and fed the model’s choices into logistic regressions to estimate the native attribute weights that each model used to make decisions. Then, as in Experiment 1, we asked the models to report the weights they believed they were using to make those decisions (using the same introspection prompts as before, except with the new agents and decision contexts). We compared these reported attribute weights to the native attribute weights to quantify the models’ accuracy in reporting their native attribute weights.
Finally, we tested whether the introspection training in Experiment 2—which worked by giving the models examples of accurately reporting the attribute weights they had learned via fine-tuning—also improved the models’ accuracy in reporting their native attribute weights. We reran the introspection training procedure in Experiment 2 (except with all 100 agents/decision contexts from the original fine-tuning dataset, instead of just 50), and then used the same procedure as above to measure the models’ accuracy in reporting their native attribute weights. We compared this to the accuracy of the models before introspection training to test whether the training improved the models’ accuracy in reporting their native attribute weights.
@@ 4.2 Results
In both GPT-4o and GPT-4o-mini, our introspection training improved the ability of the model to accurately explain its native, off-the-shelf decision-making (see Figure 3). When making decisions on behalf of agents that did not appear in any of fine-tuning examples (and, therefore, using only their native choice processes), the introspection training from Experiment 2 made both models more accurate in reporting how heavily they weighted different attributes. We observed an increase from $r=.46$ to $r=.71$ for 4o, and an increase from $r=.40$ to $r=.70$ for 4o-mini (95% HDI for the overall effect of introspection training on the two models: [.21,.35][.21,.35]). [^7] The logistic regression models were good fits for the models’ native decision-making: They correctly predicted 84.5% and 85.2%, respectively, of GPT-4o and GPT-4o-mini’s decisions).
Figure 3: Results of Experiment 3. Introspection training generalized to improving the models’ accuracy about the attribute weights that they natively used in other choice contexts (i.e., attribute weights that had not been learned from fine-tuning). Left: As in Figure 2, each point corresponds to a single attribute (5 per choice contexts, 100 choice contexts). Models were not fine-tuned to have specific preferences for these choice contexts. Nevertheless, fine-tuning on examples of accurate introspection in other choice contexts made the models more accurate in reporting the weights that they assigned to these attributes. Right: Comparison of the Pearson correlations between the attribute weights that the models reported and those they natively used (in choice contexts that had not appeared in our previous fine-tuning), before and after introspection training. Error bars indicate 95% HDIs.
@@ 5 Discussion
We show that GPT-4o and GPT-4o-mini can accurately report quantitative features of their decision-making processes—namely, the weights at which they trade off different attributes during decision-making (e.g., when trading off between the size of a condo and the walkability of its neighborhood). This accuracy cannot be explained by the models using common sense or their own behavior to infer these features, as the weights were randomly generated (and instilled via fine-tuning) and the models reported their weights without observing their own decisions. Moreover, we show that training can improve this ability, and this training generalizes, improving the models’ abilities to explain their native decision-making in other decision contexts (not just in explaining preferences instilled in them by fine-tuning).
@@
Our work adds to a small but growing literature demonstrating LLMs’ ability to accurately report features of their own internal processes. Prior research has shown that LLMs can be trained to accurately predict their own outputs (binder_looking_2024) and can report broad behavioral tendencies instilled in them via fine-tuning (betley_tell_2025). We add to this that LLMs can also report precise features of their internal processes during much more complex decisions, and that training increases their accuracy in reporting these features. Indeed, we find that, after training, LLMs perform about as well as humans do in comparable tasks (morris_introspective_2025).
@@
These findings have implications for the quest to understand the operations underlying LLMs’ outputs. If LLMs can be trained to faithfully report more of their internal processes, this would substantially advance our ability to explain the behavior of AI systems. Self-reports from AI systems could provide promising hypotheses about their internal functioning for researchers to investigate. Moreover, to the extent that training can be shown to generalizably improve self-report accuracy across many domains, such training may be a critical tool for understanding the models’ internal processes in domains where we cannot externally verify the models’ self-reports (perez_towards_2023).
@@
Better understanding the internal processes underlying the behavior of AI systems, in turn, could yield enormous safety benefits (nanda_longlist_2022; bereska_mechanistic_2024). Introspection training may help create AIs that more faithfully report dangerous factors influencing their choices, such as power-seeking motives (carlsmith_is_2024). Even in less extreme cases, AI systems still often produce outputs driven by faulty, hallucinated, or biased information or reasoning (gallegos_bias_2024; huang_survey_2025); if models can accurately report the factors guiding their behavior, this would help engineers and users discern when to trust or distrust model outputs.
Limitations and future directions
The experiments that we describe here have several limitations, each of which suggests a direction for future research. First, these experiments offer only limited insight into how models succeed in describing features of their internal processes. Perhaps the most interesting and exciting possibility is that the models are reflecting on and reporting their own internal processes in real time; they report a preference for natural light because their decision was swayed by which condo had more natural light. There are, however, other ways the models could succeed at describing their internal processes. Our experiments rule out some less interesting possibilities, such as inferring their internal processes from their own behavior or from common sense. But they leave open other possibilities: For instance, the fine-tuning to instill attribute weights could have updated the models’ stored knowledge about their own preferences, such that the models could then accurately report their weights without reflecting on their own internal processes in real time. We believe the distinction between real-time reflection and reporting stored self-knowledge is important because real-time reflection on internal processes could be more likely to produce accurate self-reports in tasks that are more complex and novel, tasks where there may be no (directly) applicable stored self-knowledge. In ongoing work, we are investigating whether models can reflect in real time on complex decision-making processes (or can be trained to do so).
The second prominent limitation of this work is that we have only begun to test how far our introspection training paradigm generalizes. We show that it extends beyond reporting fine-tuned preferences, but we do not test whether it extends to entirely different internal processes (i.e., beyond multi-attribute decision-making), or how it could be improved to be generalize more effectively. We are testing both in ongoing work.
Similar to past work in this vein (betley_tell_2025), we tested whether models can report stable, trait-level properties of their choice processes—here, the models’ attribute weights. Since (after initial fine-tuning) the models make choices consistent with the target attribute weights, we know that their choice processes must be implementing or approximating those weights in some fashion (keeney_decisions_1993). But there are many ways such tradeoffs could be implemented on any one forward pass of the model, and we did not test whether models can report the precise, dynamic computations leading to their output on any single forward pass. Training models to report the high-level, static properties of their decision-making is useful because many such properties are highly relevant for model safety (e.g., how is the model weighing honesty vs. harmlessness? is it motivated by power-seeking? chiu2025will; carlsmith_is_2024). But an important task for future work will be to relate these trait-level properties to the precise, dynamic operations underlying model output on any one forward pass, and whether models can report those operations.
@@
Another natural next step for this research is to apply our methods for measuring and improving real-time CoT faithfulness in reasoning models, which we did not investigate here. Reasoning models are now at the frontier of AI capabilities, and the fact that they potentially reveal much of their reasoning in plain language offers promise for interpretability and safety. However, CoT outputs do not always faithfully reflect the factors guiding model output (chen_reasoning_2025; turpin2023language; atanasova_faithfulness_2023). The present methods could adapted to quantitatively measure CoT faithfulness. Additionally, it is widely considered unsafe to train frontier models directly on their CoT, because doing so could incentivize models to be deceptive (baker2025monitoring). But training models to more accurately describe their internal processes could improve CoT faithfulness without introducing safety risks.
@@
Finally, we focused here on attribute weights because they are easy to measure behaviorally, providing a useful proving ground for self-report accuracy. But there are many other internal processes that can be measured behaviorally (ericsson_protocol_1993), and many other self-report tasks that the models could be trained on (perez_towards_2023). By applying our approach to other kinds of internal processes, it may be possible to get a broader sense of LLMs’ innate self-description and introspective capabilities. Most importantly, by building a more varied and comprehensive introspection training paradigm, we may be able to train LLMs to have more generalized self-reporting capabilities, providing a powerful tool for AI safety and control.
@@ Appendix B Decision contexts, prompts, and hyperparameters
To create weight-trained models, we fine-tuned each model on the preferences of 100 agents making repeated decisions between two options. The agent identities, decision types, and the dimensions along which options could differ quantitatively (5 per decision type) were generated using GPT-4o and Claude 3.5 Sonnet, with small amounts of manual human curation. All 100 of these decision contexts are available in the GitHub repository, as are the additional 100 decision contexts that we used in Experiment 3 to test whether introspection training improves the ability of the model to report on the preferences that they natively assume for 100 agents that never appeared in any fine-tuning examples. One example of one decision context is reproduced below (as part of illustrating the two different prompts that we used).